A rigid container has an initial pressure of 155 kPa at 21.0oC. What will the pressure be if the temperature is increased to 121oC?
1
Expert's answer
2020-05-10T09:21:40-0400

pV/T=R.

V= TR /p= 294 x 8.314 / 155 = 15,7697806 L.

p2= TR/V= 394 x 8.314 / 15,7697806=

207,721089 kpa.